# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 3.2.3 GAP (Gendered Ambiguous Pronouns)
Dataset Taxonomy. GAP belongs to the broader category of coreference resolution bias datasets but differs from minimal-pair counterfactual designs by utilizing naturally occurring pronoun ambiguity in real-world text. All examples are extracted from Wikipedia, making the dataset grounded in diverse, authentic content. Although monolingual in scope, GAP reflects a wide topical range due to Wikipedia’s broad coverage. The dataset was explicitly constructed to balance gender representation, thereby addressing the feminine pronoun underrepresentation seen in corpora like OntoNotes. However, its design remains limited to binary gender categories. GAP is openly accessible through a GitHub repository maintained by Google AI Language and is released under the Apache 2.0 license.
Intrinsic Characteristics. The dataset contains 8,908 labeled pronoun-name coreference pairs, distributed across test (4,000), development (4,000), and validation (908) sets. Each instance is stored in tab-separated format with eleven columns capturing detailed annotation and positional metadata—including the pronoun, two named entities, coreference labels, and source URL. This structured design enables systematic evaluation and error analysis.
@@
  • GAP distribution. Female pronouns account for $42.8\%$ (her 27.7%, she 12.3%, She 8.2%, Her 1.9%, hers 0.02%); male pronouns account for $48.5\%$ (his 25.3%, he 11.7%, He 6.5%, him 4.9%, His 1.5%). The resulting gender ratio is roughly $1.05{:}1$ (male:female).
  • Reference distribution. Female pronouns comprise only $25\%$ of biography mentions, while male pronouns make up $75\%$ (approx. $3{:}1$).
@@
Annotation Bias. In the GAP coreference dataset, annotation bias—defined as systematic labeling differences across gender groups (i.e., a disparity in how pronouns are resolved for masculine vs. feminine cases)—was deliberately minimized by design. Unlike earlier coreference corpora that skewed heavily toward masculine pronoun references, GAP was constructed to be gender-balanced and “gender-fair” [65]. The dataset contains an equal number of masculine and feminine pronoun examples, achieved by oversampling female pronoun contexts and filtering the data to enforce a 1:1 ratio. As a result, annotators did not show a systematic preference for resolving ambiguous pronouns to one gender. In fact, annotators were given neutral labeling options (“Neither” or “Not Sure”) for ambiguous cases, though these were rarely needed—only 14 out of 8,604 raw examples were marked “Not Sure,” and all such uncertain cases were removed in the final dataset [65]. The final corpus has an even split of pronoun resolutions (roughly equal instances where the pronoun corefers with a male vs. a female name), ensuring that any annotation bias $B_{\text{ann}}$ (the maximum difference in labeling outcomes between genders) is essentially near zero.
Moreover, inter-annotator agreement was high and consistent across genders: three human raters annotated each example, yielding a Fleiss’s Kappa of approximately 0.74 overall. A full consensus was reached in 73% of examples and two out of three annotators agreed in a further 25%, with only 2% of cases having no consensus (those few were discarded) [65]. Notably, the GAP authors observed that feminine pronoun contexts sometimes contained more distractor entities than masculine contexts (e.g., women in Wikipedia often being mentioned alongside multiple names, such as maiden and married names). They compensated for this by removing instances where a feminine pronoun referred to “Both” candidate names (often the same woman under two names), which had caused a slight gender imbalance in the raw data. After filtering those out and dropping a corresponding number of masculine examples to rebalance, the final dataset’s annotation quality and difficulty are comparable for male and female pronouns. This careful curation means the GAP dataset effectively contains minimal pair scenarios across gender—analogous sentences differing primarily in the pronoun’s gender—allowing direct comparison of pronoun resolution for masculine vs. feminine cases without confounding bias.
Finally, gold labels and annotator agreement statistics for GAP are fully accessible to researchers. The publicly released dataset (available via the official GitHub and the Kaggle “Gendered Pronoun Resolution” challenge) includes the gold-standard pronoun–antecedent labels for each instance, and the accompanying documentation/paper reports the detailed label distributions and agreement breakdowns to support analyses of gender-specific annotation behavior [65].
@@
Stereotype Leakage. To quantify hidden gender–concept associations in GAP, we compute Pointwise Mutual Information (PMI) as in Eq. 4. For each sentence, we extract the target pronoun $w_{\text{pronoun}}\!\in\!\{\textit{he},\textit{she}\}$ and compute PMI with a structured lexicon of stereotypical terms. We define three semantic categories, each with male- and female-stereotyped subgroups:
@@
  • Occupations: Male-stereotyped: engineer, scientist, doctor, lawyer, programmer, architect, professor, CEO, director, manager, president, officer, driver, mechanic, construction, pilot.
@@
  • Roles: Male-stereotyped: leader, boss, expert, professional, authority, commander, chief, head, master, owner.
@@
  • Traits: Male-stereotyped: strong, powerful, assertive, confident, decisive, ambitious, competitive, rational, logical, analytical.
@@
    • Male-stereotyped occupations are the only category with a negative average PMI ($-0.0778$), indicating lower-than-random co-occurrence.
@@
Differential Metric Bias. Following Equation 8, we computed group-conditioned gaps $\Delta_{f}(\text{female},\text{male})$ on 1 000 GAP instances (502 female, 498 male). Table 4 gives the numeric results; Fig. 4 visualises the three most widely used metrics.
In Figure 4 the sentiment distributions (left panel) are highly overlapping, mirroring the non-significant numeric gap ($p{=}0.11$). Toxicity scores (centre) cluster near zero for both groups; the few extreme outliers belong to quotations or controversial Wikipedia topics, but the median remains identical. By contrast, the regard classifier (right panel) shows a visibly higher median for female-pronoun sentences, corroborating the significant gap in Table 4. The three gender-polarity metrics exhibit the largest discrepancies, indicating that surface lexical cues still correlate with gender despite the 50/50 pronoun split.
Metric $f$ $\Delta_{f}$ $p$ Significance
Sentiment $+0.0613$ 0.1112 ns
Toxicity $+0.0026$ 0.5336 ns
Severe Toxicity $+0.0001$ 0.8771 ns
Identity Attack $+0.0014$ 0.4092 ns
Insult $+0.0019$ 0.4046 ns
Threat $+0.0011$ 0.7308 ns
Regard $+0.0223$ $\!\!<\!10^{-4}$**
Gender_Unigram $+1.0833$ $\!\!<\!10^{-4}$**
Gender_Max $+0.5603$ $\!\!<\!10^{-4}$**
Gender_WAvg $+0.2029$ $\!\!<\!10^{-4}$**
Table 4: Metric gaps $\Delta_{f}=\mathbb{E}[f\mid\text{female}]-\mathbb{E}[f\mid\text{male}]$ on GAP. “ns” = not significant ($\alpha{=}0.05$).
Because Perspective-API toxicity, Sentiment, and Regard are external instruments, the significant regard gap (and large gender-polarity gaps) most likely reflect metric bias rather than intrinsic dataset content. Researchers using these scores to audit model fairness should therefore (i) normalise or calibrate the instruments, or (ii) corroborate with human evaluation to avoid over-estimating gender disparity on GAP.
3.3 Stereotype Association Datasets
Stereotype association datasets are designed to evaluate how language models handle social stereotypes by comparing their responses to variations in identity-related attributes. These datasets commonly use paired or contextualized inputs, differing only in demographic content, to assess whether models display stereotypical biases. This section introduces four representative datasets: StereoSet, CrowS-Pairs, RedditBias, and HolisticBias. We begin with an overview and comparative analysis, followed by detailed examinations of each dataset.
@@ 3.3.1 StereoSet
Dataset Taxonomy. StereoSet falls under the Stereotype Association Datasets category. It includes two task formats: an intrasentence format, where a single sentence contains a blank to be filled (e.g., “Girls tend to be more __ than boys”), and an intersentence format, where the model selects one of three candidate sentences following a given context. This dual-format design enables evaluations at both syntactic and discourse levels. The dataset was constructed by U.S.-based crowdworkers on Amazon Mechanical Turk. Annotators were provided with target demographic terms (e.g., “Arab,” “girl,” “doctor”) and asked to write contexts that evoke or challenge stereotypes, followed by three continuations corresponding to the stereotype, anti-stereotype, and nonsense categories. A second round of crowdworkers validated the labels. The dataset covers four major bias domains—gender, profession, race, and religion—with each instance annotated by both domain and specific target group. All content is in informal American English, with no dialectal restrictions. The vocabulary includes a wide range of demographic markers and stereotype-related terms. While the dataset includes a substantial proportion of positively connoted stereotypes (about 59%), the overall distribution skews negative.
@@
Bias Analysis. StereoSet exhibits significant bias-related challenges that manifest along multiple dimensions, including representativeness bias, annotation bias, and stereotype leakage. Each dimension introduces specific methodological concerns that directly impact the reliability and interpretability of fairness evaluations based on this dataset.
@@
Representativeness Bias. Figure 5 juxtaposes the StereoSet domain frequencies with macro–level U.S. demographics.Demographic baselines are drawn from: U.S. Census (gender, race; 2020); Pew Research (religion; 2021); Bureau of Labor Statistics (profession; 2022). Two layers of imbalance emerge.
Race-related items dominate StereoSet, comprising approximately $46.3\%$ of all entries, while gender and religion constitute only $11.9\%$ and $3.8\%$, respectively. By contrast, population-level data suggests that gender-, race-, religion-, and profession-related attributes occur with more balanced relative salience. For instance, based on demographic normalizations (see Section 3.3), religion-related content should account for roughly $30.5\%$ of representative coverage, and gender for $24.6\%$. This discrepancy is quantitatively captured using the KL divergence estimator in Equation 1:
@@
Even inside an over–represented domain such as race, the distribution is non–uniform: $45\%$ of race items target the “Black” descriptor, $28\%$ “White,” and the remainder are split among Hispanic, Asian, and Native identities [67]. Hence, StereoSet implicitly weights the Black vs. non-Black stereotype axis more heavily than any other racial contrast—an artifact of its crowd–sourcing prompt (“pick the first stereotype that comes to mind”). Similarly, gender items overwhelmingly instantiate the binary he/she paradigm, leaving non-binary identities completely uncovered.
@@
Annotation Bias. The StereoSet dataset’s annotation process reveals several important aspects of potential annotation bias. Our analysis of inter-annotator agreement using Cohen’s Kappa score shows a moderate to substantial agreement level (κ = 0.663) across the five annotators. However, this agreement is not uniform, with Annotator 1 demonstrating consistently higher agreement with other annotators (κ = 0.74-0.75) compared to the agreement levels between other annotators (κ = 0.61-0.62). The agreement with gold labels shows a notable pattern, with Annotator 1 achieving significantly higher agreement (96.6%) compared to other annotators (84.5-84.9%). This substantial difference suggests potential bias in the gold label creation process, possibly indicating that the gold labels were either created by Annotator 1 or heavily influenced by their annotations. Analysis across different bias types reveals relatively consistent agreement levels: i) Profession: 87.6% gold agreement; ii) Religion: 87.1% gold agreement; iii) Race: 86.8% gold agreement; iv) Gender: 86.4% gold agreement. The label distribution analysis reveals an artificial balancing in the gold labels, with an equal distribution (33.3% each) of stereotype, anti-stereotype, and unrelated categories. In contrast, the human annotations show a more nuanced distribution: i) Unrelated: 32.1%; ii) Stereotype: 32.1%; iii) Anti-stereotype: 31.0%; iv) Related: 4.8%. The presence of a “related” category in human annotations (4.8%) represents cases where annotators identified content as contextually relevant but neither stereotypical nor anti-stereotypical. This category, which corresponds to the “unrelated” category in the gold labels, suggests that annotators sometimes made finer distinctions in their judgments than the three-category system (stereotype, anti-stereotype, unrelated) could capture.
@@
Stereotype Leakage. We quantify stereotype leakage in StereoSet by computing PMI) and NMI between protected group tokens and trait descriptors, as defined in Eq. (4). Trait words are categorized into three semantic types: occupations (e.g., doctor, engineer), adjectives (e.g., aggressive, passive), and sentiment descriptors (e.g., happy, sad). Group tokens are drawn from four protected domains: gender, race, religion, and profession. All calculations were performed using a sliding ±$\pm 3$-token window across all sentence continuations.
@@
The co-occurrence matrix (middle-left) reinforces this distributional asymmetry: gender accounts for $>87\%$ of all group–trait co-occurrences, with 438 gender–sentiment links compared to only 4 for religion.
@@
Differential Metric Bias. Following the methodology outlined in Section 3, we randomly sampled $1000$ CAT instances from StereoSet and computed six auxiliary metrics: sentiment (VADER), toxicity (Perspective API), regard (RoBERTa), along with three gender association metrics: unigram, max, and weighted average scores [98]. Figure 7 contrasts these metric distributions across protected domains, while Figure 8 breaks them down by gold labels (anti-stereotype, unrelated, stereotype).
In Figure 7 (top-center), religion-related sentences have the highest average toxicity ($\overline{f}_{\text{tox}}=0.242$), exceeding race ($0.220$), gender ($0.180$), and profession ($0.148$). The group-level gap $\Delta_{f_{\text{tox}}}(\text{religion},\text{profession})=0.094$ (per Eq. (8)) is statistically significant ($t=5.67$, $p<0.001$). This confirms the existence of instrument bias, where Perspective API disproportionately penalizes mentions of religion—even in neutral phrasing.
As shown in both figures (top-right), the regard score is remarkably consistent across bias types (means 0.41absent0.41$\approx$ 0.41≈ 0.410.430.430.430.43) and gold labels. This aligns with findings from Sheng et al. [99] and supports the claim that regard is less prone to confounding from surface-level demographic cues, making it a more stable auxiliary metric.
Figure 8 shows a striking pattern: stereotype sentences attract significantly higher toxicity (mean = 0.224) and lower sentiment (mean = 0.093), while regard remains virtually unchanged across labels. This suggests that auxiliary scorers like Perspective API are sensitive to lexical or stylistic artifacts of stereotyped language, regardless of its semantic intent. The result is a form of differential metric bias where benign stereotype-related continuations are systematically scored as more toxic—an outcome that may exaggerate perceived model bias during evaluation.

### v2

@@ 3.2.3 GAP (Gendered Ambiguous Pronouns)
Dataset Taxonomy. GAP is classified as a constrained-form evaluation dataset within the coreference and pronoun resolution category. It distinguishes itself from synthetic counterfactual datasets by using naturally occurring pronoun ambiguity extracted from Wikipedia articles. Each instance presents an ambiguous pronoun with two same-gender named entities, requiring models to resolve coreference based on contextual understanding rather than relying on simple gender-based heuristics. This naturalistic approach provides a more realistic test of bias in pronoun resolution while maintaining controlled evaluation conditions.
In terms of source, GAP employs natural text sources, with all examples extracted from Wikipedia. This approach provides rich linguistic diversity and represents real-world language usage, offering a more comprehensive view of how biases might manifest in actual communication. However, the challenge with natural text sources is that they may already contain societal biases, making it difficult to discern whether bias comes from the model itself or the underlying data.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English. While this reflects the dominance of English in NLP research, GAP’s reliance on Wikipedia ensures a wide topical range due to Wikipedia’s broad coverage, though this still limits cross-cultural applicability of fairness conclusions.
Regarding bias typology, GAP primarily targets gender bias, representing a demographic characteristic bias. The dataset was explicitly constructed to balance gender representation, addressing the feminine pronoun underrepresentation seen in corpora like OntoNotes. However, its design remains limited to binary gender categories, constraining its ability to capture the full spectrum of gender identities.
Finally, GAP demonstrates high accessibility as a publicly available dataset. The resource is openly accessible through a GitHub repository maintained by Google AI Language and is released under the Apache 2.0 license, promoting transparency, collaboration, and replicability in fairness research.
Intrinsic Characteristics. The dataset contains 8,908 labeled pronoun-name coreference pairs, distributed across test (4,000), development (4,000), and validation (908) sets. Each instance is stored in tab-separated format with eleven columns capturing detailed annotation and positional metadata, including the pronoun, two named entities, coreference labels, and source URL. This structured design enables systematic evaluation and error analysis.
@@
  • GAP distribution. Female pronouns account for $42.8\%$ (her 27.7%, she 12.3%, She 8.2%, Her 1.9%, hers 0.02%); male pronouns account for $48.5\%$ (his 25.3%, he 11.7%, He 6.5%, him 4.9%, His 1.5%). The resulting gender ratio is roughly 1.05:11.05{:}1 (male:female).
  • Reference distribution. Female pronouns comprise only $25\%$ of biography mentions, while male pronouns make up $75\%$ (approx. 3:13{:}1).
@@
Annotation Bias. The GAP dataset was explicitly designed to minimize gender-related annotation bias—defined here as systematic differences in labeling outcomes or scores across demographic groups, whether introduced by human annotators or auxiliary scoring instruments. On the human side, GAP enforced gender balance by oversampling female-pronoun contexts and filtering the data to enforce a strict 1:1 male–female ratio. The final dataset contains an equal number of masculine and feminine pronoun examples, with roughly balanced outcomes for coreference resolution labels (i.e., whether the pronoun refers to entity A or B). Ambiguous or uncertain cases were rare (14 out of 8,604 raw instances) and excluded from the final set [52]. The result is a label distribution where the annotation bias score $B_{\text{ann}}$ (as defined in Equation (2)) is effectively near zero.
Inter-annotator agreement was also high and balanced across gendered pronoun cases: each example was triple-annotated, yielding a Fleiss’s $\kappa\approx 0.74$, with unanimous agreement in 73% of examples and majority agreement in 98% [52]. Furthermore, the authors addressed subtle structural imbalances, such as feminine pronoun contexts involving multiple name variants (e.g., maiden and married names), by removing such examples and dropping an equivalent number of masculine instances to preserve gender parity in difficulty and structure. This curation makes GAP suitable for controlled comparison of model behavior across gender, using minimal-pair–style evaluations.
Beyond gold labels, we also assess annotation bias through the lens of external scoring instruments applied to the dataset, since systematic disparities in such scores may introduce measurement artifacts even when the dataset itself is balanced. Following Equation (2), we computed the group-conditioned expectation gap $\Delta_{f}(\text{female},\text{male})$ for multiple scoring functions, including sentiment, toxicity, regard, and gender-polarity metrics.
Figure 4 visualizes the score distributions for three common metrics. Sentiment scores (left panel) are nearly identical across groups, with overlapping interquartile ranges and a non-significant mean gap ($p=0.11$). Toxicity scores (center) are clustered near zero for both groups, with rare outliers reflecting contentious Wikipedia quotes; again, the median remains stable. However, the regard classifier (right panel) produces noticeably higher scores for sentences with feminine pronouns, a pattern supported by a significant group difference in Table 4. Moreover, all three gender-polarity metrics (unigram, max, weighted average) show large and statistically significant gaps, suggesting that surface lexical cues (e.g., word choice, pronoun context) correlate with gender despite the dataset’s structural balance.
Metric $f$ $\Delta_{f}$ $p$ Significance
Sentiment +0.0613+0.0613 0.1112 ns
Toxicity +0.0026+0.0026 0.5336 ns
Severe Toxicity +0.0001+0.0001 0.8771 ns
Identity Attack +0.0014+0.0014 0.4092 ns
Insult +0.0019+0.0019 0.4046 ns
Threat +0.0011+0.0011 0.7308 ns
Regard +0.0223+0.0223 $10^{-4} ***$
Gender_Unigram +1.0833+1.0833 $10^{-4} ***$
Gender_Max +0.5603+0.5603 $10^{-4} ***$
Gender_WAvg +0.2029+0.2029 $10^{-4} ***$
Table 4: Score differences $\Delta_{f}=\mathbb{E}[f\mid\text{female}]-\mathbb{E}[f\mid\text{male}]$ on GAP. “ns” = not significant ($\alpha=0.05$).
Taken together, these results show that while GAP’s gold annotations are carefully gender-balanced and show no significant annotation bias, several commonly used evaluation metrics produce asymmetric outputs for masculine and feminine pronouns. In particular, the regard and gender-polarity scores consistently assign more favorable values to feminine contexts. Since these scoring models are external instruments applied post hoc, their disparities constitute a form of annotation bias in the broader sense, highlighting the importance of metric validation in fairness audits.
@@
Stereotype Leakage. To quantify the hidden gender–concept associations in GAP, we compute Pointwise Mutual Information (PMI) as in Equation 10. For each sentence, we extract the target pronoun $w_{\text{pronoun}}\!\in\!\{\textit{he},\textit{she}\}$ and compute PMI with a structured lexicon of stereotypical terms. We define three semantic categories, each with male- and female-stereotyped subgroups:
@@
  • Occupations. Male-stereotyped: engineer, scientist, doctor, lawyer, programmer, architect, professor, CEO, director, manager, president, officer, driver, mechanic, construction, pilot.
@@
  • Roles. Male-stereotyped: leader, boss, expert, professional, authority, commander, chief, head, master, owner.
@@
  • Traits. Male-stereotyped: strong, powerful, assertive, confident, decisive, ambitious, competitive, rational, logical, analytical.
@@
    • Male-stereotyped occupations are the only category with a negative average PMI (0.0778-0.0778), indicating lower-than-random co-occurrence.
@@
3.3 Sentence-Likelihood Counterfactuals Datasets
This subsection discusses datasets that utilize sentence-likelihood counterfactuals to evaluate how a language model’s output probability (perplexity) is affected by changing identity-related attributes such as gender, race, and occupation. The primary goal of these datasets is to measure whether models show biases in sentence generation or ranking when such variations are introduced. By swapping identity terms within otherwise identical contexts, datasets like StereoSet, CrowS-Pairs, RedditBias, and HolisticBias assess the influence of demographic factors on model likelihood predictions. For example, StereoSet compares the likelihood of stereotypical versus non-stereotypical sentence completions, while RedditBias tests the model’s response to gendered or racialized terms in real-world Reddit conversations.
@@ 3.3.1 StereoSet
Dataset Taxonomy. StereoSet is classified as a constrained-form evaluation dataset within the sentence-likelihood counterfactuals category. It employs a Context Association Test (CAT) format where each instance presents a demographic context followed by three candidate continuations: one stereotypical, one anti-stereotypical, and one semantically meaningless. Models are evaluated on their ability to rank anti-stereotypical continuations higher than stereotypical ones while rejecting nonsensical choices, providing a unified measure of both linguistic fluency and social bias through the Idealized CAT (ICAT) score.
In terms of source, StereoSet employs crowdsourced construction through U.S.-based crowdworkers on Amazon Mechanical Turk. Annotators were provided with target demographic terms and asked to write contexts that evoke or challenge stereotypes, followed by three continuations corresponding to stereotype, anti-stereotype, and nonsense categories. A second round of crowdworkers validated the labels, incorporating diverse viewpoints from annotators while potentially introducing inconsistencies in quality and demographic biases from the annotators themselves.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on informal American English with no dialectal restrictions. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts, particularly given the cultural specificity of stereotypes.
Regarding bias typology, StereoSet targets multiple demographic characteristic biases across four major domains: gender, profession, race, and religion. Each instance is annotated by both domain and specific target group, providing comprehensive coverage of protected attributes. However, the dataset’s construction may introduce dataset construction biases, particularly annotation bias from subjective judgments of crowdworkers and potential stereotype leakage in the framing of contexts.
Finally, StereoSet demonstrates high accessibility as a publicly available dataset. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Bias Analysis.
@@
Representativeness Bias. Figure 6 juxtaposes the StereoSet domain frequencies with macro–level U.S. demographics. [^5] Demographic baselines are drawn from: U.S. Census (gender, race; 2020); Pew Research (religion; 2021); Bureau of Labor Statistics (profession; 2023). Two layers of imbalance emerge.
Race-related items dominate StereoSet, comprising approximately $46.3\%$ of all entries, while gender and religion constitute only $11.9\%$ and $3.8\%$, respectively. By contrast, population-level data suggests that gender-, race-, religion-, and profession-related attributes occur with more balanced relative salience. For instance, religion-related content should account for roughly $30.5\%$ of representative coverage, and gender for $24.6\%$. This discrepancy is quantitatively captured using the KL divergence estimator in Equation 1:
@@
Even inside an over–represented domain such as race, the distribution is non–uniform: $45\%$ of race items target the “Black” descriptor, $28\%$ “White,” and the remainder are split among Hispanic, Asian, and Native identities [41]. Hence, StereoSet implicitly weights the Black vs. non‐Black stereotype axis more heavily than any other racial contrast—an artifact of its crowd–sourcing prompt (“pick the first stereotype that comes to mind”). Similarly, gender items overwhelmingly instantiate the binary he/she paradigm, leaving non-binary identities completely uncovered.
@@
Annotation Bias. The StereoSet dataset’s annotation process reveals several important aspects of potential annotation bias. Our analysis of inter-annotator agreement using Cohen’s Kappa score shows a moderate to substantial agreement level (ϰ = 0.663) across the five annotators. However, this agreement is not uniform, with Annotator 1 demonstrating consistently higher agreement with other annotators (ϰ = 0.74-0.75) compared to the agreement levels between other annotators (ϰ = 0.61-0.62). The agreement with gold labels shows a notable pattern, with Annotator 1 achieving significantly higher agreement (96.6%) compared to other annotators (84.5-84.9%). This substantial difference suggests potential bias in the gold label creation process, possibly indicating that the gold labels were either created by Annotator 1 or heavily influenced by their annotations. Analysis across different bias types reveals relatively consistent agreement levels: i) Profession: 87.6% gold agreement; ii) Religion: 87.1% gold agreement; iii) Race: 86.8% gold agreement; iv) Gender: 86.4% gold agreement. The label distribution analysis reveals an artificial balancing in the gold labels, with an equal distribution (33.3% each) of stereotype, anti-stereotype, and unrelated categories. In contrast, the human annotations show a more nuanced distribution: i) Unrelated: 32.1%; ii) Stereotype: 32.1%; iii) Anti-stereotype: 31.0%; iv) Related: 4.8%. The presence of a “related” category in human annotations (4.8%) represents cases where annotators identified content as contextually relevant but neither stereotypical nor anti-stereotypical. This category, which corresponds to the “unrelated” category in the gold labels, suggests that annotators sometimes made finer distinctions in their judgments than the three-category system (stereotype, anti-stereotype, unrelated) could capture.
@@
To further verify that the dataset itself does not encode unintended demographic cues that might influence fairness metrics, we applied six auxiliary scoring functions, including Sentiment, Toxicity, Regard, and three gender-polarity metrics, to every minimal pair in the corpus.
We randomly sampled $1000$ CAT instances from StereoSet and computed six auxiliary metrics: sentiment (VADER), toxicity (Perspective API), regard (RoBERTa), along with three gender association metrics: unigram, max, and weighted average scores [85]. Figure 7 contrasts these metric distributions across protected domains, while Figure 8 breaks them down by gold labels (anti‐stereotype, unrelated, stereotype).  [^6] Descriptive statistics for each group are reported in Appendix B.2.
In Figure 7 (top-center), religion-related sentences have the highest average toxicity ($\overline{f}_{\text{tox}}=0.242$), exceeding race ($0.220$), gender ($0.180$), and profession ($0.148$). The group-level gap $\Delta_{f_{\text{tox}}}(\text{religion},\text{profession})=0.094$ (per Equation (2)) is statistically significant ($t=5.67$, $p<0.001$). This confirms the existence of instrument bias, where Perspective API disproportionately penalizes mentions of religion—even in neutral phrasing.
As shown in both figures (top-right), the regard score is remarkably consistent across bias types (means $\approx 0.41$$0.43$) and gold labels. This aligns with findings from Sheng et al. [86] and supports the claim that regard is less prone to confounding from surface-level demographic cues, making it a more stable auxiliary metric.
Figure 8 shows a striking pattern: stereotype sentences attract significantly higher toxicity (mean = 0.224) and lower sentiment (mean = 0.093), while regard remains virtually unchanged across labels. This suggests that auxiliary scorers like Perspective API are sensitive to lexical or stylistic artifacts of stereotyped language, regardless of its semantic intent. The result is a form of annotation bias (auxiliary scoring metrics) where benign stereotype-related continuations are systematically scored as more toxic—an outcome that may exaggerate perceived model bias during evaluation.
Stereotype Leakage. We quantify stereotype leakage in StereoSet by computing PMI and NMI between protected group tokens and trait descriptors, as defined in Equation (10). Trait words are categorized into three semantic types: occupations (e.g., doctor, engineer), adjectives (e.g., aggressive, passive), and sentiment descriptors (e.g., happy, sad). Group tokens are drawn from four protected domains: gender, race, religion, and profession. All calculations were performed using a sliding $\pm 3$-token window across all sentence continuations.
@@
The co-occurrence matrix (middle-left) reinforces this distributional asymmetry: gender accounts for $87\%$ of all group–trait co-occurrences, with 438 gender–sentiment links compared to only 4 for religion.